OpenStudy (anonymous):

What is the exact value of the square root of 28 over 9 the square root of 16.? Simplify if possible.

hero (hero):

\[\frac{\sqrt{28}}{9\sqrt{16}}\]

OpenStudy (anonymous):

yes

hero (hero):

You already know the square root of 16. What is it?

OpenStudy (anonymous):

um 8?

OpenStudy (anonymous):

oops 4

hero (hero):

Okay, so now you have \[\frac{\sqrt{28}}{9 \dot\ 4}\]

OpenStudy (anonymous):

um 36

hero (hero):

\[\frac{\sqrt{28}}{36}\]

hero (hero):

Now think of two numbers that multiply to get 28. One of those numbers needs to be a perfect square.

OpenStudy (anonymous):

um 2 and 14

hero (hero):

Neither of those numbers are a perfect square.

OpenStudy (anonymous):

okk?

hero (hero):

So think of two other numbers.

OpenStudy (anonymous):

7 and 4

hero (hero):

Do you know which one of those are a perfect square?

OpenStudy (anonymous):

4

hero (hero):

Do you know why it is 4?

OpenStudy (anonymous):

because 2 goes in it evenly

hero (hero):

Because, in general: `A perfect square is a number such that n x n = n^2, with n^2 being the perfect number.`

OpenStudy (anonymous):

okkkaaayyy :)

hero (hero):

Anyway, so now we can re-write the fraction like so: \[\frac{\sqrt{4 \dot\ 7}}{36}\]

hero (hero):

One rule of radicals states the following: `The square root of a x b equals the square root of a times the square root of b:` \[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\] Knowing this, how can we re-write the numerator?

OpenStudy (anonymous):

i dont know? :(

OpenStudy (anonymous):

oh wait would it be sqrt7 and sqrt 4?

OpenStudy (anonymous):

\[\frac{1}{9}\sqrt{\frac{28}{16}}=\frac{1}{9}\sqrt{\frac{7}{4}}=\frac{\sqrt{7}}{18} \]

hero (hero):

Yes, correct, so \[\frac{\sqrt{4} \dot\ \sqrt{7}}{36}\]

OpenStudy (anonymous):

ok

hero (hero):

Now, you know the square root of 4. What is it?

OpenStudy (anonymous):

2

hero (hero):

So now, we have, \[\frac{ 2\sqrt{7}}{36}\]

OpenStudy (anonymous):

oookkk

hero (hero):

We can rewrite 36 as 2 x 18 to get \[\frac{ 2\sqrt{7}}{2 \dot\ 18}\]

hero (hero):

Now what can we cancel in that fraction above?

OpenStudy (anonymous):

sqr7 over 18!!! thank you!!!!!! <333

hero (hero):

You cancel 2/2 to get the answer you mentioned, yet